OpenStudy (osanseviero):

How to demostrate the associative property of adittion? (for complex numbers)

OpenStudy (amistre64):

sin(a+b) = sin(a)cos(b) + sin(b)cos(a) cos(a+b) = cos(a)cos(b) - sin(a)sin(b) might be useful

OpenStudy (osanseviero):

Isn't there a simpler way?

OpenStudy (amistre64):

hmm, can we use the vector properties of complex numbers? then address that vector addition is associative?

OpenStudy (amistre64):

a complex number (a+bi) can be expressed as a vector (a,b)

OpenStudy (amistre64):

take 3 vectors, u,v,w (u+v) + w and show that you can convert it to u + (v+w)

OpenStudy (osanseviero):

Ok...That is what I was thinking to do

OpenStudy (osanseviero):

Do I put them as points or vectors, then?

OpenStudy (amistre64):

a + bi + c + di -------- a+c + (b+d)i a+c + (b+d)i m + ni ------------- (a+c+m) + (b+d+n)i or some other convoluted torture lol

OpenStudy (amistre64):

vectors is best, since its a pretty basic notion to show that vector addition is associative

OpenStudy (osanseviero):

So...what? I do the same drawing two times?

OpenStudy (amistre64):

http://www.rootmath.org/linear-algebra/algebraic-properties-of-vectors about 3:30 starts the associative proofing of vectors

OpenStudy (osanseviero):

I am trying commutative for myself, but thanks :) this will help

OpenStudy (amistre64):

yep, good luck

OpenStudy (osanseviero):

thanks

OpenStudy (osanseviero):

How to demostrate that z+(-z)=0 ?